n a farming community, 30% of the farmers grow oranges only, 10% grow lemons only and 4% grow both oranges and lemons. 1.1) What proportion of farmers in the community grow either oranges or lemons? [3 marks] 1.2) If a farmer is chosen randomly from these in the community, what is the probability that he grows neither oranges nor lemons? [marks] 1.3) Of all the farmers who grow oranges, what proportion grow lemons also?


1
Expert's answer
2022-04-21T15:14:23-0400

Let "A-" the event that a farmer grows oranges only, "P(A) =0.3;"

"B-" the event that a farmer grows lemons only, "P(B) =0.1;"

"C-" the event that a farmer grows both oranges and lemons, "P(C) =0.04."


1.1) "D-" the event that a farmer grows either oranges or lemons, "D=A\cup B."

Events "A, B" are mutually exclusive events,

"P(D) =P(A\cup B) =P(A) +P(B) =\\
=0.30+0.10=0.40=40\%."


1.2) "E-" the event that a farmer grows neither oranges nor lemons.

"F-" the event that a farmer grows oranges or lemons or both, "F=A\cup B\cup C;"

Events "A, B, C" are mutually exclusive events,

"P(F) =P(A\cup B\cup C) =\\=P(A) +P(B) +P(C) =\\
=0.30+0.10+0.04=0.44=44\%."


Events "E" and "F" are complementary events,

"E=\bar F, P(E) =1-P(F)=1-0.44=0.56=56\%."


1.3)

"\cfrac{P(C) } {P(A)+P(C) } =\cfrac{0.04} {0.3+0.04} =0.1176=11.76\%."